# On $t$-structures adjacent and orthogonal to weight structures

### v1

@@ 5.3 Some adjacent weight structures in the saturated case
Now we proceed towards extending Theorem 5.2.2(2a) to a rather nice statement on saturated categories. To formulate it in a rather general form we will define some new types of $t$-structures.
@@
Let us now relate these notions to semi-orthogonal decompositions.
@@
I.1. Then the subcategory $\underline{C}_{t=-\infty}$ is triangulated and $t$ restricts to them.
2. The restriction $t_{+}$ of $t$ to $\underline{C}_{+}$ is bounded. Moreover, $t_{+}$ is essentially bounded if $t$ is.
@@
2. We construct this weight structure as an extension of the weight structure $w_{+}$ provided by Theorem 5.2.2(I.1). As we have noted in Remark 4.2.3(2), applying Corollary 4.2.2(1) to the category $\underline{C}^{op}$ one can obtain a semi-orthogonal decomposition $(\underline{C}_{+},\underline{C}_{0})$ of $\underline{C}$ (see Remark 1.2.7(2)) with $\operatorname{Obj}\underline{C}_{0}={}^{\perp}(\operatorname{Obj}\underline{C}_{+})$. Then Proposition 3.2(5) of [BoV19] implies that $w=(\operatorname{Obj}\underline{C}_{0}\star\underline{C}_{+,w_{+\leq 0}},\underline{C}_{t\geq 0})$ is a weight structure. Here we apply Proposition 5.3.2(II) and duality to obtain the existence of the adjoint functor mentioned in loc. cit. and recall that $\underline{C}_{t\geq 0}=\underline{C}_{+,w_{+\geq 0}}$.

### v2

@@ 5.3 Some adjacent weight structures in the saturated case
Now we proceed towards extending Theorem 5.2.2(I.2a) to a rather nice statement on saturated categories. To formulate it in a rather general form we need some definitions related to $t$-structures.
@@
I.1. Then the subcategory $\underline{C}_{t=-\infty}$ is triangulated and $t$ restricts to it.
2. The restriction $t_{+}$ of $t$ to $\underline{C}_{+}$ (see Theorem 5.2.2(I.1)) is bounded below. Moreover, $t_{+}$ is essentially bounded if $t$ is.
@@
2. We construct this weight structure as an extension of the weight structure $w_{+}$ provided by Theorem 5.2.2(I.1).
As we have noted in Remark 4.2.3(2), applying Corollary 4.2.2(1) to the category $\underline{C}^{op}$ (and its semi-orthogonal decomposition $(\underline{C}_{+},\underline{C}_{t=-\infty})$; see Proposition 5.3.2(2) and Remark 1.2.7(2)) one obtains a semi-orthogonal decomposition $(\underline{C}_{+},\underline{C}_{0})$ of $\underline{C}$ with $\operatorname{Obj}\underline{C}_{0}={}^{\perp}(\operatorname{Obj}\underline{C}_{+})$. Then Proposition 3.2(5) of [BoV19] implies that $w=(\operatorname{Obj}\underline{C}_{0}\star\underline{C}_{+,w_{+\leq 0}},\underline{C}_{t\geq 0})$ is a weight structure. Here we apply Proposition 5.3.2(II) and duality to obtain the existence of the adjoint functor mentioned in loc. cit. and recall that $\underline{C}_{t\geq 0}=\underline{C}_{+,w_{+\geq 0}}$.
@@
6 Some additional remarks
In this section we make some (not really important) remarks. Most of them relate the results above to various statements in the literature. We also discuss some examples.
To relate our Definition 1.3.1(1) (of orthogonality between $w$ in $t$) to the more general definition given in [Bon10b] we recall Definition 2.5.1(1) of ibid.
Definition 6.1.
Remark 6.2.
  1. 1.
    Assume $\underline{C},\underline{C}^{\prime}\subset\underline{D}$. Then the restriction of the bi-functor $\underline{D}(-,-)$ to $\underline{C}^{op}\times\underline{C}^{\prime}$ is easily seen to give a duality $\underline{C}^{op}\times\underline{C}^{\prime}\to\operatorname{Ab}$; see Proposition 2.5.6(1) of [Bon10b].
  2. 2.
    In Definition 2.5.1(3) of [Bon10b] and §5.2 of [Bon19] orthogonality was defined with respect to an arbitrary duality $\Phi$ between $\underline{C}$ and $\underline{C}^{\prime}$; see Definition 6.1.
    Consequently, all the properties of orthogonal weight and $t$-structures established in ibid. can be applied in the setting of Definition 1.3.1(1).
Remark 6.3.
1. This paper is a certain modification of the preprint [Bon19]. One of the main distinctions is that the somewhat ad hoc (and restrictive) notion of reflection of categories (see Definition 2.2.1 of ibid.) was avoided. Respectively, Corollary 2.4.4 is a certain substitute for Theorem 2.2.5 of [Bon19]. Though the author suspects that the results of the current text are not sufficient to prove loc. cit. itself, they can be successfully applied to treat all the examples of loc. cit. known to the author; cf. Proposition 3.3.5.
Moreover, several new general results were added; this includes the theory of coproductive extensions, (most of) §4, §5.3, and subcategories corresponding to certain support conditions (see Propositions 2.3.5, 3.3.4, 3.3.5, and 4.2.6(2), and Corollary 4.2.2(2)).
2. Since [Bon19] (as well as the earlier preprint [Bon16]) is cited in some recent papers, it makes sense to note the following: Theorem 3.2.3(I) of [Bon19] (as well as Theorem 3.1.2 of [Bon16] that has been cited by several authors) coincides with Theorem 0.1 of the current paper, Theorem 3.3.1 of [Bon19] almost coincides with Theorem 3.3.2 below, Theorem 5.3.1(I.1) of ibid. is contained in Proposition 5.2.1, Theorem 5.3.1(IV) of ibid. was generalized to Theorem 5.3.3, and strongly ${\aleph_{0}}$–well generated weight structures (that were treated in §3.3 of [Bon19]) are discussed in §3.3 (yet see Remark 6.6 below).
Remark 6.4.
The methods of the current paper are not the most general among the existing methods for constructing $t$-structures. In particular, if $\underline{C}$ is a well generated triangulated category (categories of this sort are well-known to be very common) then the recent Theorem 2.3 of [Nee21] gives all those $t$-structures that are generated by sets of objects of $\underline{C}$ in the sense of Definition 3.2.1(5).
Now, if $\underline{C}$ is well generated then is generated by a set of its objects (as its own localizing subcategory; see Definition 3.2.1(3)). Thus for any smashing weight structure $w$ on it Proposition 2.3.2(10) of [Bon21] essentially says the following: there exists a set $\mathcal{P}\subset\underline{C}_{w\geq 0}$ such that the class $\underline{C}_{w\geq 0}$ is the smallest cocomplete pre-aisle that is generated by $\mathcal{P}$ in the sense of [Nee21, §0] (cf. Discussion 1.16 of ibid.). Hence Theorem 2.3 of [Nee21] implies that there exists a $t$-structure on $\underline{C}$ such that $\underline{C}_{t\geq 0}=\underline{C}_{w\geq 0}$ (cf. Remark 1.1.4(2) below). Thus loc. cit. generalizes the existence of $t$ part of our Theorem 0.1.
Remark 6.5.
Let us discuss examples to Theorem 3.2.3; cf. Remark 3.2.4.
1. First we apply the theorem to semi-orthogonal decompositions.
If $w$ is a semi-orthogonal decomposition of $\underline{C}$ (where $\underline{C}$ satisfies the Brown representability property) that is smashing as a weight structure then for the corresponding $t^{l}$ the couple $(\underline{C}_{t^{l}\leq 0},\underline{C}_{t^{l}\geq 0})$ is a semi-orthogonal decomposition as well; see Proposition 1.3.5. Note also that $(\underline{C}_{t^{l}\leq 0},\underline{C}_{t^{l}\geq 0})$ is a (cosmashing) weight structure.
Thus if we apply Proposition 1.3.5 to the semi-orthogonal decomposition $(\underline{C}_{t^{l}\leq 0},\underline{C}_{t^{l}\geq 0})$ we obtain that the embedding $\underline{C}_{2}\to\underline{C}$ possesses a right adjoint as well; note that $\underline{C}_{t^{l}\geq 0}=\underline{C}_{w\geq 0}$. Thus $\underline{C}_{2}$ is admissible in $\underline{C}$ in the sense of [BoK89, Definition 2.5] and the embedding $\underline{C}^{\prime}\to\underline{C}$ may be completed to a gluing datum (cf. [BBD82, §1.4] or [Nee01, §9.2]).
2. Let us say a little more on smashing weight structures.
Theorem 4.5(2) of [PoS16] and Theorem 3.2.1 of [Bon22] enable one to check whether two weight structures generated by a set of compact objects are distinct.
Other notable statements related to the construction of smashing weight structures are Theorem 2.3.4(3) of [Bon22] (it says that any perfect set of objects in a smashing triangulated category generates a smashing weight structure) and Theorem 3.1.3 and 3.2.2 of [BoS19] that treat weight-exact localization functors.
Remark 6.6.
Let us make a funny observation concerning Theorem 3.3.2.
According to the well-known Lemma 4.4.5 of [Nee01], the assumptions of Theorem 3.3.2 imply that $\underline{C}^{({\aleph_{0}})}$ is the retraction-closure (see §1.1) of the category ${\underline{C}^{0}}$. Note however that $w^{0}$ does not extend to $\underline{C}^{({\aleph_{0}})}$ in general (even though it extends to $\underline{C}$!); see §3.1 of [BoS18]. [^9] In any example of this sort $w^{0}$ is neither bounded above nor below; see Theorem 2.2(II.2) of ibid. Consequently, it does makes some sense not to assume ${\underline{C}^{0}}=\underline{C}^{({\aleph_{0}})}$ in our theorem.
For the same reason, the weight structures provided by our Theorem 3.3.2 don’t have to be strongly ${\aleph_{0}}$–well generated in the sense of Remark 3.3.4(1) of [Bon22] (even if one assumes that ${\underline{C}^{0}}$ is essentially small; note that this is equivalent to the compact generation of $\underline{C}$). This observation possibly suggests that the notion of strong ${\aleph_{0}}$–well generation should be slightly generalized.